OpenStudy (anonymous):

I NEED HELP! Simplify each expression. Use positive exponents. 1. m^3n^-6p^0 2. (a^4b^-3/ab^-2) 3. (x^-2y^-4x^-3)-2

OpenStudy (haseeb96):

listen @_brooke16 on any digit if the power is zero then it will be equal to 1 okay like this 1^0= 1 and if any digit has -ve power then just reciprocal it and change the power into +ve p^-7 = 1/p^7 okay

OpenStudy (haseeb96):

try to do number one equation

OpenStudy (anonymous):

Okay let me try real quick

OpenStudy (anonymous):

I still don't get how to do it.

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

\[m^3 n^-6 p^0\]

OpenStudy (anonymous):

so this is what you have (well, the 6 should be little)

OpenStudy (anonymous):

@kah.x yes

OpenStudy (anonymous):

so you essentialy have m^3 times n^-6 times p^0

OpenStudy (anonymous):

what he was saying is any number^0 is equal to 1 1^0 = 1 2^0 = 1 3^0 = 1 so when you have a variable^0 it is equal to 1

OpenStudy (anonymous):

so you essentially have m^3 n^-6 1

OpenStudy (anonymous):

make sense?